The heat evolved from the combustion of 0.94g ethanol raises the temperature of

250 cm3 water contained in a copper beaker by 20.0 °C. [Specific heat capacity of

solution = 4.18 Jg-1 °C-1; density of solution = 1.0 g cm-3]

i) What is the enthalpy of combustion per mol of ethanol?

ii) Calculate the temperature rise when 1.50 g ethanol is used to heat 500 cm3

water in the same copper beaker.


1
Expert's answer
2021-04-25T03:50:51-0400

I) 0.94g × 20°c= 19.8

19.8/10 × 4.8

= 9.504

ii) ∆T = 500-250 = 250

250/(4.18) = 59.81

= 59.81-20

= 39.81°C